OpenStudy (anonymous):

A company distributes free pencils to all the students of x schools. Each school has (x + 3) classes. The number of students in each class is 4 more than the number of classes in each school. Each student is given (x + 1) pencils. Part A: Write an expression to show the total number of pencils distributed by the company in x schools. (4 points) Part B: What does x(x + 3) represent? When simplified, what would be the degree and classification of this expression? (4 points) Part C: How can you calculate the total number of students in each school? (2 points)

OpenStudy (anonymous):

This is my last question. Anybody out there please help me. FAN and MEDAL

OpenStudy (dbzfan836):

ok

OpenStudy (dbzfan836):

@geerky42 come now, we have a prob

geerky42 (geerky42):

well, there is x schools and each school have x+3 classes. so for total of classes we have is x(x+3). "The number of students in each class is 4 more than the number of classes in each school." So let s be number of student in each class, so we have equation s = 4+(x+3) we can combine like term and get s = x+7. so the total students is x(x+3)(x+7) do you understand what I just did? can you figure the expression for total of pencils?

OpenStudy (anonymous):

Im lost. Im sorry.

OpenStudy (dbzfan836):

@mathmale here

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

ok I am officially lost to

OpenStudy (anonymous):

but I might be able to figure it out give me a min

OpenStudy (anonymous):

@sourwing would it be 4+(x+3)(x+1)? for part a

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

I'm so lost that I am debating on whether or not to watch the show Lost. I love Jack and Kate, they rock.

OpenStudy (anonymous):

im not very helpful . sorry:( i'm supposed to know this. too many word are confusing me

OpenStudy (anonymous):

its a five question assessment so i can skip this but itll automatically drop my grade to an 80.

OpenStudy (anonymous):

im being stupid and im sorry but im not taking credit either. check this out: http://openstudy.com/study#/updates/5307eaaae4b022471a6681a4

OpenStudy (anonymous):

no u r NOT going to skip it k

OpenStudy (anonymous):

Damn are you kidding?! lmao

OpenStudy (anonymous):

Best Response Medals 1 a) x(x+3)(x+7)(x+1) = x^4+11x^3+31x^2+21x. If you're unsure of which, I'd include both. b) (x+3) = classes per school, x = schools, so x(x+3) = total classes. x(x+3)=x^2+3x, degree 2 (greatest exponent). c) Chop off the pencils per student in the original equation (drop "(x+1)"). x(x+3)(x+7) or x^3+10x^2+21x I give credit to @tomhoe this is probably the correct answer